Test of the apacite package and bibstyle

Erik Meijer

Department of Psychometrics and Research Methodology

Leiden University PO Box 9555 2300 RB Leiden The Netherlands E-Mail: Meijer@rulfsw.LeidenUniv.nl

June 7, 1999

Abstract

In this document, the examples of the APA manual (3rd ed., 1984) are cited to test the apacite package and bibstyle. A description of the apacite package and bibstyle can be found in the file

apacite.tex. See also the readme.apa file. The file apacite.sty

must be in a directory that TEX can find. The files apacite.bst and testAPA.bib must be in directories that BibTEX can find.

1 The in-text citations

The following enumerated list gives the example number, and the two basic in-text citation commands, both in a full and short form, of all examples in the APA manual.

1. (Paivio, 1975)

(Paivio, 1975) Paivio (1975) Paivio (1975)

2. (Becker & Seligman, 1981) (Becker & Seligman, 1981) Becker and Seligman (1981) Becker and Seligman (1981)

3. (Horowitz, Post, French, Wallis, & Siegelman, 1981) (Horowitz et al., 1981) Horowitz, Post, French, Wallis, and Siegelman (1981) Horowitz et al. (1981)

4. (Winston et al., 1983)

(Winston et al., 1983) Winston et al. (1983) Winston et al. (1983)

5. (Corcoran & Williamson, in press)

(Corcoran & Williamson, in press) Corcoran and Williamson (in press) Corcoran and Williamson (in press)

6. (Gardner, 1981)

(Gardner, 1981) Gardner (1981) Gardner (1981)

7. (Staff, 1980)

(Staff, 1980) Staff (1980) Staff (1980)

8. (“Study Finds”, 1982)

(“Study Finds”, 1982) “Study Finds” (1982) “Study Finds” (1982)

9. (Lublin, 1980)

(Lublin, 1980) Lublin (1980) Lublin (1980)

10. (O’Neill, 1982)

(O’Neill, 1982) O’Neill (1982) O’Neill (1982)

11. (Glaser & Bond, 1981) (Glaser & Bond, 1981) Glaser and Bond (1981) Glaser and Bond (1981)

12. (Hood & Bloom, 1979)

(Hood & Bloom, 1979) Hood and Bloom (1979) Hood and Bloom (1979)

13. (Battig & Montague, 1969)

(Battig & Montague, 1969) Battig and Montague (1969) Battig and Montague (1969)

14. (Corrigan, Dell, Lewis, & Schmidt, 1980) (Corrigan et al., 1980) Corrigan, Dell, Lewis, and Schmidt (1980) Corrigan et al. (1980)

15. (Misumi & Fujita, 1982)

(Misumi & Fujita, 1982) Misumi and Fujita (1982) Misumi and Fujita (1982)

16. (Koczkas, Holmberg, & Wedin, 1981) (Koczkas et al., 1981) Koczkas, Holmberg, and Wedin (1981) Koczkas et al. (1981)

17. (Cialdini, Petty, & Cacioppo, 1981) (Cialdini et al., 1981) Cialdini, Petty, and Cacioppo (1981) Cialdini et al. (1981)

18. (Assink & Verloop, 1977)

(Assink & Verloop, 1977) Assink and Verloop (1977) Assink and Verloop (1977)

19. (Stutte, 1972)

(Stutte, 1972) Stutte (1972) Stutte (1972)

20. (Beatty, 1982)

(Beatty, 1982) Beatty (1982) Beatty (1982)

21. (Ree, Mullins, Mathews, & Massey, 1983) (Ree et al., 1983) Ree, Mullins, Mathews, and Massey (1983) Ree et al. (1983)

22. (Strunk & White, 1979)

(Strunk & White, 1979) Strunk and White (1979) Strunk and White (1979)

23. (American Psychiatric Association, 1980) (American Psychiatric Association, 1980) American Psychiatric Association (1980) American Psychiatric Association (1980)

24. (Letheridge & Cannon, 1980)

(Letheridge & Cannon, 1980) Letheridge and Cannon (1980) Letheridge and Cannon (1980)

25. (“College Bound Seniors”, 1979)

(“College Bound Seniors”, 1979) “College Bound Seniors” (1979) “College Bound Seniors” (1979)

26. (Cohen, 1977)

(Cohen, 1977) Cohen (1977) Cohen (1977)

27. (Wilson & Fraser, 1977–1978)

(Wilson & Fraser, 1977–1978) Wilson and Fraser (1977–1978) Wilson and Fraser (1977–1978)

28. (Piaget & Inhelder, 1951)

(Piaget & Inhelder, 1951) Piaget and Inhelder (1951) Piaget and Inhelder (1951)

29. (Luria, 1965/1969)

(Luria, 1965/1969) Luria (1965/1969) Luria (1965/1969)

30. (Gurman & Kniskern, 1981)

(Gurman & Kniskern, 1981) Gurman and Kniskern (1981) Gurman and Kniskern (1981)

31. (Woodward, in press)

(Woodward, in press) Woodward (in press) Woodward (in press)

32. (Epstein, 1967)

(Epstein, 1967) Epstein (1967) Epstein (1967)

33. (Davydov, 1972)

(Davydov, 1972) Davydov (1972) Davydov (1972)

34. (Freud, 1923/1961)

(Freud, 1923/1961) Freud (1923/1961) Freud (1923/1961)

35. (Sluzki & Beavin, 1965/1977)

(Sluzki & Beavin, 1965/1977) Sluzki and Beavin (1965/1977) Sluzki and Beavin (1965/1977)

36. (National Institute of Mental Health [NIMH], 1982) (NIMH, 1982) National Institute of Mental Health [NIMH] (1982) NIMH (1982)

37. (Tandy, 1980)

(Tandy, 1980) Tandy (1980) Tandy (1980)

38. (Gottfredson, 1980)

(Gottfredson, 1980) Gottfredson (1980) Gottfredson (1980)

39. (Brush, 1979)

(Brush, 1979) Brush (1979) Brush (1979)

40. (Kandel, 1976)

(Kandel, 1976) Kandel (1976) Kandel (1976)

41. (Newport, 1975)

(Newport, 1975) Newport (1975) Newport (1975)

42. (Ben-Yishay, 1981)

(Ben-Yishay, 1981) Ben-Yishay (1981) Ben-Yishay (1981)

43. (Life Insurance Marketing and Research Association [LIMRA], 1978) (LIMRA, 1978) Life Insurance Marketing and Research Association [LIMRA] (1978) LIMRA (1978)

44. (Chaddock, Carlson, & Hamilton, 1974)

(Chaddock et al., 1974) Chaddock, Carlson, and Hamilton (1974) Chaddock et al. (1974)

45. (Thumin, Craddick, & Barclay, 1973)

(Thumin et al., 1973) Thumin, Craddick, and Barclay (1973) Thumin et al. (1973)

46. (Singh, 1980)

(Singh, 1980) Singh (1980) Singh (1980)

47. (Brener, 1979)

(Brener, 1979) Brener (1979) Brener (1979)

48. (Pendar, 1982)

(Pendar, 1982) Pendar (1982) Pendar (1982)

49. (Foster-Havercamp, 1981/1982) (Foster-Havercamp, 1981/1982) Foster-Havercamp (1981/1982) Foster-Havercamp (1981/1982)

50. (Devins, 1981)

(Devins, 1981) Devins (1981) Devins (1981)

51. (Ryerson, 1983)

(Ryerson, 1983) Ryerson (1983) Ryerson (1983)

52. (Cameron, 1981)

(Cameron, 1981) Cameron (1981) Cameron (1981)

53. (Gottfredson, 1978)

(Gottfredson, 1978) Gottfredson (1978) Gottfredson (1978)

54. (Palm, 1981)

(Palm, 1981) Palm (1981) Palm (1981)

55. (Herbert, 1983)

(Herbert, 1983) Herbert (1983) Herbert (1983)

56. (Bailey, 1980)

(Bailey, 1980) Bailey (1980) Bailey (1980)

57. (Carmody, 1982)

(Carmody, 1982) Carmody (1982) Carmody (1982)

58. (Bowers, 1982)

(Bowers, 1982) Bowers (1982) Bowers (1982)

59. (Newman, 1982)

(Newman, 1982) Newman (1982) Newman (1982)

60. (Maas & Gluck, 1979)

(Maas & Gluck, 1979) Maas and Gluck (1979) Maas and Gluck (1979)

61. (Clark, 1976)

(Clark, 1976) Clark (1976) Clark (1976)

62. (Miller, Miller, & Kline, 1975) (Miller et al., 1975) Miller, Miller, and Kline (1975) Miller et al. (1975)

63. (Fernandes, 1972)

(Fernandes, 1972) Fernandes (1972) Fernandes (1972)

References

American Psychiatric Association. (1980). Diagnostic and statistical manual of mental disorders (3rd ed.). Washington, DC: Author.

Assink, E. M. H., & Verloop, N. (1977). Het aanleren van deel-geheel relaties in het aanvankelijk rekenonderwijs [Teaching part-whole relations in elementary mathematics instruction]. Pedagogische Studiën, 54, 130– 142.

Bailey, A. C. (Ed.). (1980, May). Newsletter of the Committee on Women in Athletics. (Available from [name and address])

Battig, W. F., & Montague, W. E. (1969). Category norms of verbal items in 56 categories: A replication and extension of the Connecticut category norms. Journal of Experimental Psychology Monographs, 80 (3, Pt. 2).

Beatty, J. (1982). Task-evoked pupillary responses, processing load, and the structure of processing resources. Psychological Bulletin, 91, 276–292.

Becker, L. J., & Seligman, C. (1981). Welcome to the energy crisis. Journal of Social Issues, 37 (2), 1–7.

Ben-Yishay, Y. (Ed.). (1981). Working approaches to remediation of cognitive deficits in brain damaged persons (Rehabilitation Monograph No. 62). New York: New York University Medical Center, Institute of Rehabilitation Medicine.

Bowers, K. S. (1982). Deeper into “Deeper into hypnosis” [Review of Deeper into hypnosis]. Contemporary Psychology, 27, 223–224.

Brener, J. (1979, October). Energy, information, and the control of heart rate. Paper presented at the meeting of the Society for Psychophysiological Research, Cincinnati, OH.

Brush, L. (1979). Why women avoid the study of mathematics: A longitudinal study (Contract No. 400-77-0099). Washington, DC: National Institute of Education.

Cameron, S. E. (1981). Educational level as a predictor of success. Unpublished manuscript.

Carmody, T. P. (1982). A new look at medicine from the social perspective [Review of Social contexts of health, illness, and patient care]. Contemporary Psychology, 27, 208–209.

Chaddock, T. E., Carlson, G. M., & Hamilton, C. L. (1974). Gastric emptying of a nutritionally balanced liquid diet in the rhesus monkey. In E. E. Daniel (Ed.), Proceedings of the Fourth International Symposium on Gastrointestinal Motility (pp. 83–92). Vancouver, British Columbia, Canada: Mitchell Press.

Cialdini, R. B., Petty, R. E., & Cacioppo, J. T. (1981). Attitude and attitude change. Annual Review of Psychology, 32, 357–404.

Clark, K. B. (Speaker). (1976). Problems of freedom and behavior modification (Cassette Recording No. 7612). Washington, DC: American Psychological Association.

Cohen, J. (1977). Statistical power analysis for the behavioral sciences (rev. ed.). New York: Academic Press.

College bound seniors. (1979). Princeton, NJ: College Board Publications.

Corcoran, D. L., & Williamson, E. M. (in press). Unlearning learned helplessness. Journal of Personality and Social Psychology.

Corrigan, J. D., Dell, D. M., Lewis, K. N., & Schmidt, L. D. (1980). Counseling as a social influence process: A review [Monograph]. Journal of Counseling Psychology, 27, 395–441.

Davydov, V. V. (1972). De introductie vam het begrip grootheid in de eerste klas van de basisschool: Een experimenteel onderzoek [The introduction of the concept of quantity in the first grade of the primary school: An experimental study]. In C. F. Van Parreren & J. A. M. Carpay (Eds.), Sovjetpsychologen aan het woord (pp. 227–289). Groningen, The Netherlands: Wolters-Noordhoff.

Devins, G. M. (1981). Helplessness, depression, and mood in endstage renal disease. Unpublished doctoral dissertation, McGill University, Montreal.

Epstein, A. N. (1967). Oropharyngeal factors in feeding and drinking. In C. F. Code (Ed.), Handbook of physiology: Sec. 6. Alimentary canal: Vol. 1. Control of food and water intake (pp. 197–218). Bethesda, MD: American Physiological Society.

Fernandes, F. D. (1972). Theoretical prediction of interference loading on aircraft stores: Part 1. Subsonic speeds [Computer program]. Pomona, CA: General Dynamics, Electro Dynamics Division. (National Aeronautics and Space Administration Report No. NASA CR-112065-1; Acquisition No. LAR-11249)

Foster-Havercamp, M. E. (1982). An analysis of the relationship between preservice teacher training and directed teaching performance (Doctoral dissertation, University of Chicago, 1981). Dissertation Abstracts International, 42, 4409A.

Freud, S. (1961). The ego and the id. In J. Strachey (Ed. and Trans.), The standard edition of the complete psychological works of Sigmund Freud (Vol. 19, pp. 3–66). London: Hogarth Press. (Original work published 1923)

Gardner, H. (1981, December). Do babies sing a universal song? Psychology Today, pp. 70–76.

Glaser, R., & Bond, L. (Eds.). (1981). Testing: Concepts, policy, practice, and research [Special issue]. American Psychologist, 36 (10).

Gottfredson, G. D. (1978). Why don’t vocational interests predict job satisfaction better than they do? Unpublished manuscript, Johns Hopkins University, Center for Social Organization of Schools, Baltimore.

Gottfredson, L. S. (1980). How valid are occupational reinforcer pattern scores? (Report No. CSOS-R-292). Baltimore, MD: Johns Hopkins University, Center for Social Organization of Schools. (ERIC Document Reproduction Service No. ED 182 465)

Gurman, A. S., & Kniskern, D. P. (1981). Family therapy outcome research: Knowns and unknowns. In A. S. Gurman & D. P. Kniskern (Eds.), Handbook of family therapy (pp. 742–775). New York: Brunner/Mazel.

Herbert, C. A. (1983). [Facilitating Learning Efficiency: Assessment Scores]. Unpublished raw data.

Hood, L., & Bloom, L. (1979). What, when, and how about why: A longitudinal study of early expressions of causality. Monographs of the Society for Research in Child Development, 44 (6, Serial No. 181).

Horowitz, L. M., Post, D. L., French, R. S., Wallis, K. D., & Siegelman, E. Y. (1981). The prototype as a construct in abnormal psychology: 2. Clarifying disagreement in psychiatric judgments. Journal of Abnormal Psychology, 90, 575–585.

Kandel, D. (1976). Study of high school students: Student questionnaire, Wave 1, Fall 1971. In A. Nehemkis, M. A. Macari, & D. J. Lettieri (Eds.), Drug abuse instrument handbook (Research Issues No. 12, pp. 259–260). Rockville, MD: National Institute of Education.

Koczkas, S., Holmberg, G., & Wedin, L. (1981). A pilot study of the effect of the 5-HT uptake inhibitor, zimelidine, on phobic anxiety. Acta Psychiatrica Scandinavica, 63 (Suppl. 290), 328–341.

Letheridge, S., & Cannon, C. R. (Eds.). (1980). Bilingual education: Teaching English as a second language. New York: Praeger.

Life Insurance Marketing and Research Association. (1978). Profits and the AIB in United States ordinary companies (Research Rep. No. 1978-6). Hartford, CT: Author.

Lublin, J. S. (1980, December 5). On idle: The unemployed shun much mundane work, at least for a while. The Wall Street Journal, pp. 1, 25.

Luria, A. R. (1969). The mind of a mnemonist (L. Solotaroff, Trans.). New York: Avon Books. (Original work published 1965)

Maas, J. B. (Producer), & Gluck, D. H. (Director). (1979). Deeper into hypnosis [Film]. Englewood Cliffs, NJ: Prentice-Hall.

Miller, W., Miller, A., & Kline, G. (1975). The CPS 1974 American national election study [Machine-readable data file]. Ann Arbor: University of Michigan, Center for Political Studies (Producer). Ann Arbor: Inter-University Consortium for Political and Social Research (Distributor).

Misumi, J., & Fujita, M. (1982). Effects of PM organizational development in supermarket organization. Japanese Journal of Experimental Social Psychology, 21, 93–111. (From Psychological Abstracts, 1982, 68, Abstract No. 11474)

National Institute of Mental Health. (1982). Television and behavior: Ten years of scientific progress and implications for the eighties (DHHS Publication No. ADM 82-1195). Washington, DC: U.S. Government Printing Office.

Newman, P. (1982, January). [Interview with William Epstein, editor of JEP: Human Perception and Performance]. APA Monitor, pp. 7, 39.

Newport, E. L. (1975). Motherese: The speech of mothers to young children (Tech. Rep. No. 52). San Diego: University of California, Center for Human Information Processing.

O’Neill, G. W. (1982, January). In support of DSM-III [Letter to the editor]. APA Monitor, p. 4.

Paivio, A. (1975). Perceptual comparisons through the mind’s eye. Memory & Cognition, 3, 635–647.

Palm, E. J. (1981). Associative learning and recognition. Manuscript submitted for publication.

Pendar, J. E. (1982). Undergraduate psychology majors: Factors influencing decisions about college, curriculum and career. Dissertation Abstracts International, 42, 4370A–4371A. (University Microfilms No. 82-06, 181)

Piaget, J., & Inhelder, B. (1951). La genèse de l’idée de hasard chez l’enfant [The origin of the idea of danger in the child]. Paris: Presses Universitaires de France.

Ree, M. J., Mullins, C. J., Mathews, J. J., & Massey, R. H. (1983). Vocational aptitude battery: Item and factor analyses. Psychological Documents, 13, 3. (Ms. No. 2469)

Ryerson, J. F. (1983). Effective management training: Two models. Unpublished master’s thesis, Clarkson College of Technology, Potsdam, NY.

Singh, R. (1980, February). Multiplying versus differential-weight averaging as integration rule in attribution of gift size. In C. M. Bhatia, (Chair), Dimensions of information processing. Symposium conducted at the meeting of the Indian Science Congress Association, Calcutta.

Sluzki, C. E., & Beavin, J. (1977). Symmetry and complementarity: An operational definition and a typology of dyads. In P. Watzlawick & J. H. Weakland (Eds. and Trans.), The interactional view (pp. 71–87). New York: Norton. (Reprinted from Acta Psiquiátrica y Psicológica de America Latina, 1965, 11, 321–330)

Staff. (1980, September 1). Professionals face tax rises as IRS targets personal-service corporations. Behavior Today, p. 5.

Strunk, W., Jr., & White, E. B. (1979). The elements of style (3rd ed.). New York: Macmillan.

Study finds free care used more. (1982, April). APA Monitor, p. 14.

Stutte, H. (1972). Transcultural child psychiatry. Acta Paedopsychiatrica, 38 (9), 229–231.

Tandy, S. (1980). Development of behavioral techniques to control hyperaggressiveness in young children (CYC Report No. 80-3562). Washington, DC: Council on Young Children. (NTIS No. P880-143282)

Thumin, F. J., Craddick, R. A., & Barclay, A. G. (1973). Meaning and compatibility of a proposed corporate name and symbol [Summary]. Proceedings of the 81st Annual Convention of the American Psychological Association, 8, 835–836.

Wilson, J. G., & Fraser, F. C. (Eds.). (1977–1978). Handbook of teratology (Vols. 1–4). New York: Plenum Press.

Winston, B. L., Reinhart, M. L., Sacker, J. R., Gottlieb, W., Oscar, B. B., & Harris, D. P. (1983). Effect of intertrial delays on retardation of learning. Journal of Experimental Psychology: Animal Behavior Processes, 9, 581–593.

Woodward, J. T. (in press). Children’s learning systems. In J. T. Woodward, A. Pimm, S. S. Keenan, M. N. Blum, H. A. Hamner, & P. Sellzner (Eds.), Research in cognitive development: Vol. 1. Local cognition in children. New York: Springer.